# Knowledge Distillation with Refined Logits

### v2

@@ 5.1 Settings
Type Teacher ResNet32$\times$4 ResNet32$\times$4 ResNet32$\times$4 WRN-40-2 WRN-40-2 VGG13 ResNet50
79.42 79.42 79.42 75.61 75.61 74.64 79.34
Student SHN-V2 WRN-16-2 WRN-40-2 ResNet8$\times$4 MN-V2 MN-V2 MN-V2
71.82 73.26 75.61 72.50 64.60 64.60 64.60
Feature FitNet 73.54 74.70 77.69 74.61 68.64 64.16 63.16
AT 72.73 73.91 77.43 74.11 60.78 59.40 58.58
RKD 73.21 74.86 77.82 75.26 69.27 64.52 64.43
CRD 75.65 75.65 78.15 75.24 70.28 69.73 69.11
OFD 76.82 76.17 79.25 74.36 69.92 69.48 69.04
ReviewKD 77.78 76.11 78.96 74.34 71.28 70.37 69.89
SimKD 78.39 77.17 79.29 75.29 70.10 69.44 69.97
CAT-KD 78.41 76.97 78.59 75.38 70.24 69.13 71.36
Logit KD 74.45 74.90 77.70 73.97 68.36 67.37 67.35
CTKD 75.37 74.57 77.66 74.61 68.34 68.50 68.67
DKD 77.07 75.70 78.46 75.56 69.28 69.71 70.35
LA 75.14 74.68 77.39 73.88 68.57 68.09 68.85
RC 75.61 75.17 77.58 75.22 68.72 68.66 68.98
LR 76.27 76.10 78.73 75.26 69.02 69.78 70.38
RLD (ours) 77.56 76.14 78.91 76.12 69.75 69.97 70.76
Table 2: Top-1 accuracy (%) on the CIFAR-100 validation set when the teacher and student models are heterogeneous. The best and second best results of logit distillation are highlighted in bold and underlined text, respectively. For the case where the best result of feature distillation is better than the best result of logit distillation, we highlight it with italic text. The reported results are the mean of three trials.
@@
We conduct the experiments on two standard image classification datasets: CIFAR-100 (Krizhevsky, Hinton et al. 2009) and ImageNet (Russakovsky et al. 2015). CIFAR-100 comprises 100 distinct classes, with a total of 50,000 images in the training set and 10,000 images in the validation set. Each image in this dataset is of the size 32$\times$32 pixels. ImageNet presents a larger and more complex dataset, encompassing 1,000 classes. It includes 1.28 million images in the training set and 50,000 images in the validation set, with each image resolution being 224$\times$224 pixels after pre-processing.
@@
Models used by teachers and students include ResNet (He et al. 2016), WideResNet (WRN) (Zagoruyko and Komodakis 2016), VGG (Simonyan and Zisserman 2015), ShuffleNet (SHN) (Ma et al. 2018; Zhang et al. 2018), and MobileNet (MN) (Howard et al. 2017; Sandler et al. 2018). The experimental results contain both the distillation of heterogeneous and homogeneous teacher-student models.
@@
The compared methods involved in the experiments include feature distillation and logit distillation methods. Feature distillation methods include FitNet (Romero et al. 2014), AT (Komodakis and Zagoruyko 2017), RKD (Park et al. 2019), CRD (Tian, Krishnan, and Isola 2020), OFD (Heo et al. 2019a), ReviewKD (Chen et al. 2021), SimKD (Chen et al. 2022), and CAT-KD (Guo et al. 2023). Logit distillation methods include KD (Hinton, Vinyals, and Dean 2015), CTKD (Li et al. 2023), DKD (Zhao et al. 2022), LA (Wen, Lai, and Qian 2021), RC (Cao et al. 2023), and LR (Lan et al. 2024). Notably, the latter three methods are based on correction techniques and are re-implemented by us in a unified training framework.
We follow the conventional experimental settings of previous works (Tian, Krishnan, and Isola 2020; Zhao et al. 2022; Sun et al. 2024a). For more details on implementation, please refer to the appendix.
@@ 5.2 Main Results
Type Teacher ResNet32$\times$4 VGG13 WRN-40-2 WRN-40-2 ResNet56 ResNet110 ResNet110
79.42 74.64 75.61 75.61 72.34 74.31 74.31
Student ResNet8$\times$4 VGG8 WRN-40-1 WRN-16-2 ResNet20 ResNet32 ResNet20
72.50 70.36 71.98 73.26 69.06 71.14 69.06
Feature FitNet 73.50 71.02 72.24 73.58 69.21 71.06 68.99
AT 73.44 71.43 72.77 74.08 70.55 72.31 70.65
RKD 71.90 71.48 72.22 73.35 69.61 71.82 69.25
CRD 75.51 73.94 74.14 75.48 71.16 73.48 71.46
OFD 74.95 73.95 74.33 75.24 70.98 73.23 71.29
ReviewKD 75.63 74.84 75.09 76.12 71.89 73.89 71.34
SimKD 78.08 74.89 74.53 75.53 71.05 73.92 71.06
CAT-KD 76.91 74.65 74.82 75.60 71.62 73.62 71.37
Logit KD 73.33 72.98 73.54 74.92 70.66 73.08 70.67
CTKD 73.39 73.52 73.93 75.45 71.19 73.52 70.99
DKD 76.32 74.68 74.81 76.24 71.97 74.11 71.06
LA 73.46 73.51 73.75 74.98 71.24 73.39 70.86
RC 74.68 73.37 74.07 75.43 71.63 73.44 71.41
LR 76.06 74.66 74.42 75.62 70.74 73.52 70.61
RLD (ours) 76.64 74.93 74.88 76.02 72.00 74.02 71.67
Table 3: Top-1 accuracy (%) on the CIFAR-100 validation set when the teacher and student models are homogeneous. The same convention is used as in Table 2.
@@
The top-1 validation accuracy (%) comparison results of RLD and other distillation approaches are reported in Table 2 (heterogeneous distillation pairs) and Table 3 (homogeneous distillation pairs). We can see that RLD is either the optimal or suboptimal logit distillation algorithm in all cases, and is optimal in most cases. This underscores the superiority of RLD and accentuates the significance of making corrections to teacher predictions. Concurrently, while feature distillation sometimes outperforms logit distillation, the optimal algorithm for feature distillation demonstrates greater instability, failing to perform well consistently with a certain algorithm. Moreover, feature distillation incurs a longer training duration and necessitates a more intricate algorithm design, potentially impeding its practical applicability.
@@
The top-1 and top-5 validation accuracy (%) comparison results of RLD and other distillation approaches are reported in Table 4. On this more challenging dataset, RLD successfully outperforms all existing feature and logit distillation algorithms, consistently achieving optimal performance and demonstrating the superiority of our approach.
5.3 Visualizations
In this section, we use visualization to more intuitively represent the differences between the different methods. The teacher model is ResNet32$\times$4, and the student model is ResNet8$\times$4.
Feature Visualization.
We employ the t-SNE (Van der Maaten and Hinton 2008) to visualize the output features of student models derived from KD and RLD. As depicted in Figure 4, the features obtained through RLD exhibit superior discriminative properties.
Figure 4: Visualized features learned by KD and RLD with t-SNE on the CIFAR-100 validation set.
Figure 5: Visualized teacher-student logit difference learned by DKD and RLD on the CIFAR-100 validation set. To better display, 100 classes are reshaped into a 10$\times$10 matrix.
Logit Difference Visualization.
We quantify the absolute difference in logits for each class between teacher and student models obtained via DKD and RLD, visualizing these results using the heat map in Figure 5. Despite RLD outperforming DKD, it is observed that the logit discrepancy yielded by RLD is larger than that of DKD. This observation aligns with our anticipation, given that RLD rectifies certain inaccuracies in teacher knowledge and provides students with greater autonomy in formulating their own predictions. This finding underscores that an unconsidered alignment with teacher knowledge may not be the optimal strategy, and we believe that correction-based approaches deserve more attention and research.
@@ 5.4 Extensions
$L_{\text{CE}}$ $L_{\text{SCD}}$ $L_{\text{MCD}}$ Accuracy
72.50
73.55
75.64
76.64
Table 5: Ablation study on the importance of each component in RLD. Top-1 accuracy (%) on the CIFAR-100 validation set is reported. The teacher is ResNet32$\times$4, and the student is ResNet8$\times$4.
Ablation Study.
We perform ablation study on the components of RLD and the results are shown in Table 5. The results demonstrate that each component of RLD effectively contributes to enhanced performance. Notably, when $L_{\text{SCD}}$ or $L_{\text{MCD}}$ are deployed separately (rows 2 and 3), the optimal results are achieved at different temperature settings. However, when integrating $L_{\text{SCD}}$ and $L_{\text{MCD}}$ (row 4), keeping the same temperature for both components outperforms setting different temperatures. This suggests that the interplay of losses is also critical to achieving better performance.
@@
Table 6: Top-1 accuracy (%) on the CIFAR-100 validation set when training with logit standardization technique. The student is MN-V2. The optimal results are highlighted in bold text. The reported results are the mean of three trials.
@@
We investigate the efficacy of each method when supplemented with logit standardization technique (Sun et al. 2024a). The results are shown in Table 6. The optimal results achieved by RLD underscore its superior performance and the vast potential of its integration with other methodologies.
@@ 6 Conclusion
Existing knowledge distillation methods did not consider the impact of incorrect teacher predictions on students, or arbitrarily corrected predictions and disrupted class correlation. In this paper, we introduce Refined Logit Distillation (RLD) to address these issues. RLD enables teacher models to impart two distinct forms of knowledge to the student models: sample confidence and masked correlation. It effectively mitigates over-fitting and eliminates potential misinformation from the teacher models, while maintaining class correlation. Consequently, it enables student models to acquire more valuable knowledge. Experimental results demonstrate the superiority of our proposed method.
@@
There are a few directions to improve our proposed RLD. For instance, dynamic temperature (Li et al. 2023) and meta-learning (Hospedales et al. 2021) techniques can be used to tune the hyper-parameters. Additional strategies such as data augmentation (Cubuk et al. 2019) and sample selection (Lan et al. 2024) can be employed to distill high-quality samples. Besides, combining RLD with state-of-the-art feature distillation methods may be a promising avenue of exploration to further improve the distillation performance. We consider extending correction-based knowledge distillation to the feature domain, utilizing techniques such as Class Activation Mapping (Wang et al. 2020).
Appendix A Training Efficiency
@@
Figure 6 presents the training time, accuracy, and extra parameters of various methods. Notably, feature distillation methods (FitNet, OFD, ReviewKD, and CRD) yield a substantial performance enhancement compared to traditional KD method, albeit at the cost of a significant increase in training time and extra parameters. In contrast, our RLD method, maintaining the same training time as KD without extra parameters introduced, delivers superior performance.
Appendix B Implementation Details
We follow the conventional experimental settings of previous works (Tian, Krishnan, and Isola 2020; Zhao et al. 2022; Sun et al. 2024a) and use Pytorch (Paszke et al. 2019) for our experiments.
B.1 CIFAR-100
When training on CIFAR-100, the batch size, epoch number, weight decay, and momentum are set to 64, 240, 5e-4, and 0.9, respectively. The initial learning rates are 0.01 for ShuffleNet and MobileNet, and 0.05 for other model architectures. The learning rate is divided by 10 at 150, 180 and 210 epochs. The optimizer is SGD (Sutskever et al. 2013). The training data is augmented using RandomCrop and RandomHorizontalFlip operators.
For the hyper-parameters involved in RLD, we follow DKD (Zhao et al. 2022) and LSKD (Sun et al. 2024a) to set different values for different distillation pairs. We always set $\alpha$ to 1, and determine the optimal $\beta$ and $\tau$ using grid search from the range of $\{2,4,8,16\}$ and $\{2,3,4,5\}$, respectively.

### v3

@@ 5.1 Settings
Type Teacher ResNet32$\times$4 ResNet32$\times$4 WRN-40-2 WRN-40-2 VGG13 ResNet50
79.42 79.42 75.61 75.61 74.64 79.34
Student SHN-V2 WRN-40-2 ResNet8$\times$4 MN-V2 MN-V2 MN-V2
71.82 75.61 72.50 64.60 64.60 64.60
Feature FitNet 73.54 77.69 74.61 68.64 64.16 63.16
AT 72.73 77.43 74.11 60.78 59.40 58.58
RKD 73.21 77.82 75.26 69.27 64.52 64.43
CRD 75.65 78.15 75.24 70.28 69.73 69.11
OFD 76.82 79.25 74.36 69.92 69.48 69.04
ReviewKD 77.78 78.96 74.34 71.28 70.37 69.89
SimKD 78.39 79.29 75.29 70.10 69.44 69.97
CAT-KD 78.41 78.59 75.38 70.24 69.13 71.36
Logit KD 74.45 77.70 73.97 68.36 67.37 67.35
CTKD 75.37 77.66 74.61 68.34 68.50 68.67
DKD 77.07 78.46 75.56 69.28 69.71 70.35
LA 75.14 77.39 73.88 68.57 68.09 68.85
RC 75.61 77.58 75.22 68.72 68.66 68.98
LR 76.27 78.73 75.26 69.02 69.78 70.38
RLD (ours) 77.56 78.91 76.12 69.75 69.97 70.76
Table 2: Top-1 accuracy (%) on the CIFAR-100 validation set when the teacher and student models are heterogeneous. The same convention is used as in Table 1.
@@
We conduct the experiments on two standard image classification datasets: CIFAR-100 [19] and ImageNet [30]. CIFAR-100 comprises 100 distinct classes, with a total of 50,000 images in the training set and 10,000 images in the validation set. Each image in this dataset is of size 32$\times$32 pixels. ImageNet presents a larger and more complex dataset, encompassing 1,000 classes. It includes 1.28 million images in the training set and 50,000 images in the validation set, with each image resolution being 224$\times$224 pixels after pre-processing.
@@
Models used by teachers and students include ResNet [9], WideResNet (WRN) [47], VGG [34], ShuffleNet (SHN) [24, 49], and MobileNet (MN) [14, 32].
@@
We emphasize that the experimental results in this paper are presented after exhaustive reading of the papers and codes of existing works, and with a focus on fair comparisons between methods. Therefore, MLKD [15] is excluded from the comparison due to differing experimental settings. The compared methods include feature distillation (FitNet [29], AT [18], RKD [26], CRD [39], OFD [10], ReviewKD [4], SimKD [3], and CAT-KD [8]) and logit distillation (KD [12], CTKD [21], DKD [50], LA [42], RC [1], and LR [20]) methods. The performance metrics of all comparison methods, except for the latter three correction-based approaches (LA, RC, and LR), are sourced directly from LSKD [35]. To ensure experimental fairness, we implemented these three correction-based approaches and our proposed RLD method following the experimental setups commonly used in prominent studies (e.g., LSKD [35], DKD [50], and CRD [39]). Additional implementation details are provided in the Appendix.
@@ 5.2 Main Results
The top-1 validation accuracy (%) comparison results of RLD and other distillation approaches are reported in Table 1 (homogeneous distillation pairs) and Table 2 (heterogeneous distillation pairs). We can see that RLD is either the optimal or suboptimal logit distillation algorithm in all cases, and is optimal in most cases. This underscores the superiority of RLD and accentuates the significance of making corrections to teacher predictions. While feature distillation can sometimes outperform logit distillation, its optimal method varies across teacher-student pairs, and its longer training time and complex algorithm design may hinder practical applicability.
@@
The top-1 and top-5 validation accuracy (%) comparison results of RLD and other distillation approaches are reported in Table 3. On this more challenging dataset, RLD successfully outperforms all existing feature and logit distillation algorithms, consistently achieving optimal performance and demonstrating its superiority.
Analysis.
Examining the experimental results detailed in Tables 2, 1 and 3, it is clear that the performance improvement brought about by RLD is more substantial on the ImageNet dataset than on the CIFAR-100 dataset. We presume that this discrepancy stems from the varying accuracy levels of the teacher models on the respective training sets. As shown in Figure 4, the teacher model exhibits high training accuracy on the CIFAR-100 dataset, while it shows comparatively lower accuracy on the ImageNet dataset. Given that RLD aligns with DKD when the teacher predictions are accurate, the high training accuracy on the CIFAR-100 dataset might hinder substantial divergence between these two methods. Conversely, on the ImageNet dataset, where the training accuracy is lower, RLD outperforms DKD by achieving a more substantial improvement.
Figure 4: Proportion of predictions from teacher models on the training set. (a) ResNet56. (b) ResNet50.
@@ 5.3 Extensions
Teacher ResNet56 ResNet110 VGG13
72.34 74.31 74.64
Student WideResNet-40-2
75.61
KD 76.72 77.37 76.86
DKD 77.34 77.70 77.45
RLD (ours) 78.03 78.28 77.88
$\Delta$ +0.69 +0.58 +0.43
Table 4: Top-1 accuracy (%) on the CIFAR-100 validation set when distilling with inferior teachers. Optimal results are highlighted in bold. The reported results are the mean of three trials.
Reversed Knowledge Distillation.
We explore a unique scenario termed reversed knowledge distillation [46], where the teacher performs worse than the student. This study investigates the feasibility of using an inferior teacher model to enhance the performance of a superior student model, particularly in situations where sourcing a more capable teacher model proves challenging. As shown in Table 4, among all distillation pairs, RLD achieves the best performance, and shows a large performance gap compared to DKD. Moreover, the accuracy difference $\Delta$ between RLD and DKD shows that poorer teacher model performance leads to greater RLD improvement over DKD. This can be attributed to two main factors: firstly, the use of the inferior teacher allows RLD to be better distinguished from DKD; secondly, the unique setup of reversed knowledge distillation imposes more stringent demands on the quality of knowledge transferred, thereby underscoring the effectiveness of RLD in refining distilled knowledge.
@@
Table 5: Top-1 accuracy (%) on the CIFAR-100 validation set when training with logit standardization technique LSKD [35]. The optimal results are highlighted in bold text. The reported results are the mean of three trials.
@@
We investigate the efficacy of each method when supplemented with logit standardization technique LSKD [35]. The results are shown in Table 5. The optimal results achieved by RLD underscore its superior performance and the vast potential of its integration with other methodologies.
Figure 5: Visualized teacher-student logit discrepancy learned by DKD and RLD on the CIFAR-100 validation set. For better visualization, 100 classes are reshaped into a 10$\times$10 matrix. The teacher is ResNet32$\times$4, and the student is ResNet8$\times$4.
Logit Discrepancy Visualization.
We calculate the mean absolute error (MAE) of logits for each class between teacher and student models obtained via DKD and RLD, visualizing these results using the heat map in Figure 5. Despite RLD outperforming DKD, it is observed that the logit discrepancy yielded by RLD is larger than that of DKD. This observation aligns with our anticipation, given that RLD rectifies certain inaccuracies in teacher knowledge and provides students with greater autonomy in formulating their own predictions. This finding underscores that an unconsidered alignment with teacher knowledge may not be the optimal strategy, and we believe that correction-based approaches deserve more attention and research.
Ablation Study.
We perform an ablation study on the components of RLD, and the results are shown in Table 6. The results demonstrate that each component of RLD effectively contributes to enhanced performance. Notably, while masking all classes with values greater than (denoted as “g”) those of the true class would similarly eliminate misinformation and preserve class correlations, this masking strategy (denoted as $M_{\text{g}}$) inadvertently integrates true class-related knowledge into both masked correlation and sample confidence. This overlap may create conflicts between the resulting losses, thereby hindering performance improvement. Figure 6 presents a toy example for illustration. After applying $M_{\text{g}}$, the distillation objective of MCD requires that the probabilities of classes B, C, and D be close. This conflicts with the objective of SCD, which enforces a probability of 0.4 for class B, since the sum of class probabilities cannot exceed 1. In contrast, this issue does not arise when using $M_{\text{ge}}$. Therefore, we opt for $M_{\text{ge}}$ as the masking strategy.
$L_{\text{CE}}$ $L_{\text{SCD}}$ $L_{\text{MCD}}$ Accuracy
$M_{\text{g}}$ $M_{\text{ge}}$
72.50
73.55
75.50
75.64
75.53
76.64
Table 6: Ablation study on the importance of each component in RLD. Top-1 accuracy (%) on the CIFAR-100 validation set is reported. The teacher is ResNet32$\times$4, and the student is ResNet8$\times$4. The reported results are the mean of three trials.
Figure 6: A toy example illustrates how the $M_{\text{g}}$ masking strategy can lead to loss conflict. In this example, the probabilities of classes B, C, and D in the teacher distribution are close, and the value corresponding to B is slightly larger than those of C and D.
Figure 7: Impact of the hyper-parameters $\alpha$ ($L_{\text{SCD}}$) and $\beta$ ($L_{\text{MCD}}$) on the CIFAR-100 validation set. By default, for both distillation pairs, $\alpha=1$ and $\beta=4$ (corresponding to the accuracies reported in Section 5.2, marked with dashed lines). The reported results are the mean of three trials.
Hyper-parameter Analysis.
We investigate the impact of the hyper-parameters $\alpha$ and $\beta$, which correspond to the importance of $L_{\text{SCD}}$ and $L_{\text{MCD}}$, respectively. As shown in Figure 7, a more detailed hyper-parameter search can significantly enhance the effectiveness of RLD. Notably, the optimal hyper-parameter configurations vary significantly across different distillation pairs, which fundamentally explains why existing studies [50, 35] do not adopt fixed hyper-parameter settings. Additionally, we reveal an important phenomenon: unlike DKD, where optimal performance is achieved at $\alpha=1$ [50], RLD generally prefers larger $\alpha$ values. This difference may stem from the following mechanism: when the teacher model makes incorrect predictions, smaller $\alpha$ values in the DKD loss can mitigate the negative impact of incorrect teacher knowledge but also limit the transfer of knowledge to some extent. In contrast, RLD refines the knowledge to effectively eliminate interference from incorrect information, allowing it to adapt to larger $\alpha$ values and further improve model performance.
@@ 6 Conclusion
Existing knowledge distillation methods do not consider the impact of incorrect teacher predictions on students. Alternatively, teacher outputs are arbitrarily corrected, disrupting class correlations. In this paper, we introduce Refined Logit Distillation (RLD) to address these issues. RLD enables teacher models to impart two distinct forms of knowledge to the student models: sample confidence and masked correlation. It effectively mitigates over-fitting and eliminates potential misinformation from the teacher models, while maintaining class correlations. Experimental results demonstrate the superiority of RLD.
@@
There are a few directions to improve our proposed RLD. For instance, dynamic temperature [21] and meta-learning [13] techniques can be used to tune the hyper-parameters. Additional strategies such as data augmentation [6] and sample selection [20] can be employed to distill high-quality samples. Besides, combining RLD with state-of-the-art feature distillation methods may be a promising avenue of exploration to further improve the distillation performance. We consider extending correction-based knowledge distillation to the feature domain, utilizing techniques such as Class Activation Mapping [41].
Acknowledgment
Wujie Sun and Can Wang are supported by the National Natural Science Foundation of China (No. 62476244), ZJU-China Unicom Digital Security Joint Laboratory and the advanced computing resources provided by the Supercomputing Center of Hangzhou City University.
Appendix A Class Correlation
We present results from a pilot experiment, affirming that scenarios similar to the one depicted in Figure 1 do indeed exist. We choose samples specifically from 3 classes in CIFAR-100: leopard, television, and dinosaur. Utilizing a well-trained model, we obtain the top 10 output logits. The corresponding results are displayed in Figure 8.
Figure 8(a), (b), and (c) illustrate that the model’s class correlations align with our understanding, regardless of whether the prediction is correct or not. For instance, leopard exhibits higher similarity to tiger, while television resembles clock more closely. Across all samples, if the label is “leopard”, the model does not improve prediction confidence or accuracy by decreasing the logit values of the label’s actually similar classes (such as “tiger”).
Figure 8(d) reveals that even in the output logits of a single sample, this inter-class relationship is well-maintained. For the sample labeled “dinosaur”, the top 3 predictions contain “leopard” and “television”. Consequently, within its top 10 predicted classes, the proportion of classes similar to “leopard” or “television” is significantly high, while classes resembling “dinosaur” are rare. In this case, simply correcting the true class (e.g., using augment or swap operations [1, 20, 42]) can result in significant deviations from the overall class correlations shown in Figure 8(c), which could negatively impact the distillation performance.
Appendix B Training Efficiency
Figure 9 presents the training time, accuracy, and extra parameters of various methods. Notably, feature distillation methods (FitNet, OFD, ReviewKD, and CRD) yield a substantial performance enhancement compared to KD, albeit at the cost of a significant increase in training time and extra parameters. In contrast, our RLD method, maintaining the same training time as KD and introducing no extra parameters, delivers superior performance.
@@ Appendix B Training Efficiency
Appendix C Implementation Details
We follow the conventional experimental settings of previous works [39, 50, 35] and use Pytorch [27] for our experiments.
C.1 CIFAR-100
When training on CIFAR-100, the batch size, number of epochs, weight decay, and momentum are set to 64, 240, 5e-4, and 0.9, respectively. The initial learning rates are 0.01 for ShuffleNet and MobileNet, and 0.05 for other model architectures. The learning rate is divided by 10 at 150, 180 and 210 epochs. The optimizer is SGD [38]. The training data is augmented using RandomCrop and RandomHorizontalFlip operators.
[…]